# Low-Degree Polynomials Are Good Extractors

### v1

@@ 5.3 Existential results for sumset extractors via a uniformly random function
where $(\mathbf{X}_{i},\mathbf{Y}_{j})$ are independent flat $k^{\prime}$-sources with $k^{\prime}=k-t=(1-\beta)\frac{k}{2}-O(1)$ and supports $\mathcal{X}_{i}$ and $\mathcal{Y}_{j}$ satisfying $E(\mathcal{X}_{i},\mathcal{Y}_{j})\leq\ell^{2}\cdot 2^{2(k-t)}=\ell^{2}\cdot 2^{2k^{\prime}}$, for $T=2^{t}$ and some $(p_{i,j})_{i,j\in[T]}$. Therefore, it suffices to show that a uniformly random function $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ satisfies $f(\mathbf{X}^{\prime}+\mathbf{Y}^{\prime})\approx_{\varepsilon}\mathbf{U}_{1}$ simultaneously for all pairs $(\mathbf{X}^{\prime},\mathbf{Y}^{\prime})$ of independent flat $k^{\prime}$-sources with $E(\mathbf{X}^{\prime},\mathbf{Y}^{\prime})\leq\ell^{2}2^{2k^{\prime}}$ with the desired probability.
@@ 6.1 Low-degree, low-error two-source extractors
In this section we show the existence of a low-error two-source extractor for arbitrary linear min-entropy $k=\delta n$ computable by a degree $\leq 4$ polynomial. In contrast, the best known explicit constructions of low-error two-source extractors require min-entropy $k\approx 0.44n$ [Bou05, Lew19]. More precisely, we have the following.
@@
For every $\delta>0$ there is $\zeta>0$ such that there exists a $(k=\delta n,\varepsilon=2^{-\zeta n})$-two-source extractor $\mathsf{Ext}:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ computable by a $2n$-variate $\mathbb{F}_{2}$-polynomial of degree $\leq 4$. Moreover, such a function $\mathsf{Ext}$ can be sampled with probability $0.99$ in time $O(n^{3})$, and $\mathsf{Ext}(x,y)$ can be evaluated in time $O(n^{3})$ for every input $(x,y)$.
We prove Theorem 8 by combining a previously known connection between subspace-evasive sets over $\mathbb{F}_{2}$ and low-error two-source extractors for linear min-entropy, and a construction of such subspace-evasive sets from degree-$2$ polynomials. This connection was originally conditional on the Approximate Duality conjecture [BR15]. However, this conjecture is now a theorem since it is implied by Polynomial Freiman-Ruzsa, which was recently proved by Gowers, Green, Manners, and Tao [GGMT23].
@@
We say that a function $f:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ has rank $r$ if there exist functions $g_{1},g_{2}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{r}$ such that
@@
Equivalently, $f$ has rank $r$ if and only if its $2^{n}\times 2^{n}$ truth table has rank $r$ over $\mathbb{F}_{2}$.
@@
Ben-Sasson, Lovett, and Ron-Zewi [BLR12] obtained an entropy-error tradeoff for this result, which allows us to drop the min-entropy $k^{\prime}$ to $O(n/\log n)$ while having error $2^{-\Omega(\sqrt{n})}$. For the sake of exposition, we do not use this alternative version, but remark that it can be applied here as well.
@@
A set $S\subseteq\mathbb{F}_{2}^{r}$ is said to be $(\ell,k)$-subspace-evasive if $|S\cap V|<2^{k}$ for any $-\ell$ subspace $V$ of $\mathbb{F}_{2}^{r}$. An injective function $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{r}$ is $(\ell,k)$-subspace-evasive if $h(\mathbb{F}_{2}^{n})$ is $(\ell,k)$-subspace-evasive.
@@
The following lemma relates the subspace-evasiveness of $h$ and the dispersion and rank of $\mathsf{Ext}_{h}$.
@@
Given an injective $(\ell=r/2,k)$-subspace-evasive function $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{r}$, it follows that $\mathsf{Ext}_{h}$ is a $k$-two-source disperser of rank at most $r$.
@@
Fix any two sets $A,B\subseteq\mathbb{F}_{2}^{n}$ of size $2^{k}$ each. We have that $\dim h(A),\dim h(B)\geq r/2+1$, since $h$ is $(r/2,k)$-subspace-evasive. In particular, this means that $\dim h(A)+\dim h(B)>r+1$. By Lemma 21, we conclude that $\mathsf{Ext}_{h}(A,B)=\langle h(A),h(B)\rangle=\{0,1\}$. ∎
@@
For every $C,\gamma,\rho>0$ there exists $\eta>0$ such that the following holds for all $n$: Let $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{r}$ be an injective $(\ell=r/2,k=\rho n)$-subspace-evasive function with $r\leq Cn$. Then, $\mathsf{Ext}_{h}(x,y)=\langle h(x),h(y)\rangle$ is a $(k^{\prime}=(\rho+\gamma)n,\varepsilon=2^{-\eta n})$-two-source extractor. Moreover, if $h$ is explicit, then so is $\mathsf{Ext}_{h}$.
@@
With Theorem 10 in mind, it remains to construct an $(\ell,k)$-subspace-evasive function $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n+r}$ with $r=O(n)$, $k=o(n)$, and $\ell=\frac{n+r}{2}$. Furthermore, we would like to construct $h$ in a way that $\mathsf{Ext}_{h}$ is computable by a low-degree $\mathbb{F}_{2}$-polynomial. Towards this, we consider the function
@@
where $f_{1},\dots,f_{r}\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ are random degree $\leq d$ polynomials.
@@
Let $k=10dn^{1/d}$ and $r=11n$. Then, with probability at least $1-2^{-3n^{2}}$ over the sampling of $f_{1},\dots,f_{r}$ we have that $h$ is an injective $(\ell=\frac{n+r}{2},k)$-subspace-evasive function.
Sudakov and Tomon [ST23] also consider subspace-evasive sets induced by collections of multivariate polynomials, but they analyze instead the setting where the field is large and both the dimension and intersection size are small.
We need the following lemma.
Lemma 23 ([BHL12, Lemma 1.4] and [KS05, Theorem 1.5], adapted.).
@@
Set $r=11n$. Sample $r$ random degree-$2$ polynomials $f_{1},\dots,f_{r}\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ and define $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n+r}$ as $h(x)=(x_{1},\dots,x_{n},f_{1}(x),\dots,f_{r}(x))$. Then, with probability at least $1-2^{-3n^{2}}$ over the sampling of $f_{1},\dots,f_{r}$ the following holds: For any $\delta>0$ there exists $c>0$ such that the function $\mathsf{Ext}_{h}:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ given by
@@
6.2 Improved impossibility results for extracting from polynomial and variety sources
@@
We show that the same construction we used to obtain low-degree subspace-evasive sets above also works for subspace-evasiveness if we use our new Lemma 14 instead of Lemma 23.
@@
There exist constants $c,C>0$ such that the following holds for all positive integers $k\geq t\geq 200d$ such that $d\geq 2$ and $\binom{t/100}{\leq d/2}\geq 4d^{2}(2e)^{d}$. Let $f_{1},\dots,f_{r}:\mathbb{F}_{2}^{k}\to\mathbb{F}_{2}$ be independent random degree $\leq d$ polynomials with
@@
which is at most $2^{-r\tau^{\prime}/2}$ provided that $r\geq\frac{4\tau k}{\tau^{\prime}-4\tau}=\frac{4k}{(2e)^{-d}d^{-2}\tau-4}$, The result follows by noting that our hypotheses implies that $(2e)^{-d}d^{-2}\tau-4\geq(2e)^{-d}d^{-2}\tau/2$, and so it suffices to choose any $r\geq 8d^{2}(2e)^{d}k/\tau$. ∎
@@
for a sufficiently large constant $C>0$. The third equality uses the fact that $k\leq n$ and $\binom{a}{\leq b}\geq\binom{a}{b}\geq\left(\frac{a}{b}\right)^{b}$. Using this and setting $t=\alpha\log n$ for a sufficiently small constant $\alpha>0$ yields the following corollary of Theorem 12, which states that there exist polynomial NOBF sources [^16] Following [CGG24, Definition 4], a degree-$d$ polynomial NOBF source $\mathbf{Z}\sim\mathbb{F}_{2}^{n}$ with min-entropy $k$ is a degree-$d$ polynomial source with the following structure: There exists a set $G\subseteq[n]$ of $k$ good coordinates which are independent and uniformly distributed, and each coordinate outside $G$ is computed by a degree $\leq d$ polynomial of the coordinates in $G$. of extremely high min-entropy that avoid even very tiny sumsets.
@@
For any constant $c>0$ there exists a constant $C>0$ such that the following holds for all positive integers $n\geq C$ and degrees $d\geq 2$. There exists a degree-$d$ polynomial NOBF source $\mathbf{Z}\sim\mathbb{F}_{2}^{n}$ with min-entropy $H_{\infty}(\mathbf{Z})\geq n-n\left(\frac{Cd}{\log n}\right)^{d/2}$ such that $\mathsf{supp}(\mathbf{Z})$ does not contain any sumset $A+B$ with $|A|,|B|\geq n^{c}$.
Corollary 7 generalizes [CGG24, Theorem 5.11], which worked for $d=2$ only, to degrees $d>2$, and it relies on a simpler proof. Since degree-$d$ polynomial NOBF sources with min-entropy $k$ are also degree-$d$ variety sources with min-entropy $k$ [CGG24, Claim 1], we obtain the following immediate corollary of Corollary 7.
Corollary 8.
There exists a constant $C>0$ such that for any integer $n\geq C$ the following holds. Sumset dispersers cannot be used to disperse from degree-$d$ polynomial NOBF sources over $\mathbb{F}_{2}^{n}$ with min-entropy $k=n-n\left(\frac{Cd}{\log n}\right)^{d/2}$ in a black-box manner. In particular, this also means that sumset dispersers cannot be used to disperse from degree-$d$ variety sources with the same min-entropy $k$ in a black-box manner.
@@
In order to be able to conclude that $k^{\prime}$-sumset disperser is also a disperser for another class of sources $\mathcal{C}$ using only its black-box property (i.e., that it is not constant on any sumset $A+B$ with $|A|,|B|\geq 2^{k^{\prime}}$), then it must be the case that for every $\mathbf{X}\in\mathcal{C}$ we have that $\mathsf{supp}(\mathbf{X})$ contains a sumset $A+B$ with $|A|,|B|\geq 2^{k^{\prime}}$. Since we can take $k^{\prime}=\alpha\log n$ for an arbitrarily small constant $\alpha>0$ and there are no sumset dispersers for such $k^{\prime}$ (because every function is constant on an $\Omega(\log n)$-dimensional subspace, and affine sources are sumset sources), the result follows. ∎

### v2

@@ 5.3 Impossibility results
We may alternatively see a function $f:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ of left-degree at most $\leq d$ as a collection of $2^{n}$ polynomials $\{f_{y}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}\}_{y\in\mathbb{F}_{2}^{n}}$, each of degree $\leq d$, where $f_{y}(x):=f(x,y)$. Without loss of generality, we assume that $f_{y}(0)=0$ for at least $2^{n-1}$ choices of $y$. The desired result follows if we can find sets $X,Y\subseteq\mathbb{F}_{2}^{n}$, each of size at least $2^{cdn^{1/d}}$, such that $f_{y}(x)=0$ for every $y\in Y$ and $x\in X$. Indeed, we can then take $\mathbf{X}$ and $\mathbf{Y}$ to be uniformly distributed over $X$ and $Y$, respectively.
Towards this end, consider sampling a subspace $V\subseteq\mathbb{F}_{2}^{n}$ of dimension
$t=\left\lfloor\frac{d}{e}\left(\frac{en}{4d^{3}}-1\right)^{1/d}\right\rfloor$
by iteratively sampling $v_{i}$ uniformly at random from $\mathbb{F}_{2}^{n}\setminus\operatorname{span}(v_{1},\dots,v_{i-1})$ for $i=1,\dots,t$. Looking ahead, we will choose $c>0$ small enough (i.e., smaller than $\frac{1}{4\log(2e)}$) so that $d\leq\frac{t-1}{2}$ for $d\leq c\log n$.
Using Lemma 16, the probability that some fixed polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ of degree at most $d$ with $f(0)=0$ satisfies $f(x)=0$ for all $x\in V$ is at least
$\prod_{i=0}^{t-1}\frac{2^{n-\sum_{j=0}^{d-1}(d-j)\binom{i}{j}}-2^{i}}{2^{n}}$ $\geq\prod_{i=0}^{t-1}\frac{2^{n-1-\sum_{j=0}^{d-1}(d-j)\binom{i}{j}}}{2^{n}}$ (16)
$=2^{-t-\sum_{i=0}^{t-1}\sum_{j=0}^{d-1}(d-j)\binom{i}{j}}$
$\geq 2^{-t-d^{2}t(et/d)^{d-1}}$ (17)
$\geq 2^{-n/2}.$ (18)
Equation 16 holds because
$n-\sum_{j=0}^{d-1}(d-j)\binom{t-1}{j}\geq n-d^{2}\left(\frac{et}{d}\right)^{d-1}\geq 3n/4\geq 2t,$
which follows from the fact that $d\leq\frac{t-1}{2}$ and by our choice of $t$. Equation 17 uses the fact that
$\sum_{i=0}^{t-1}\sum_{j=0}^{d-1}(d-j)\binom{i}{j}\leq d^{2}t\left(\frac{et}{d}\right)^{d-1}<n/4,$
again by our choice of $t$.
Now, if $\mathbf{A}\sim[2^{n}]$ denotes the number of polynomials in the collection $\{f_{y}:y\in\mathbb{F}_{2}^{n},f_{y}(0)=0\}$ such that $f_{y}(v)=0$ for all $v\in V$, then by linearity of expectation and Equation 18, we have
$\mathbb{E}[\mathbf{A}]\geq 2^{n-1}\cdot 2^{-n/2}=2^{n/2-1}\geq 2^{t}.$
Therefore, by an averaging argument, there exist a $t$-dimensional subspace $V$ and a set $Y\subseteq\mathbb{F}_{2}^{n}$ of size at least $2^{t}$ such that $f_{y}(V)\equiv 0$ for all $y\in Y$. Taking $X=V$ and this $Y$ (and $\mathbf{X}$ and $\mathbf{Y}$ uniformly distributed over these subsets, respectively) concludes the argument. ∎
@@ 6.1 Low-degree, low-error two-source extractors
Recall that in Theorem 2, we showed that a random degree $d$ polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is an $\varepsilon$-extractor for $n$-bit sumset sources with min-entropy $k=O(d(n/\varepsilon^{2})^{1/\lfloor d/2\rfloor})$. While this can handle any range of parameters, notice that in order to achieve “low error” $\varepsilon=k^{-\omega(1)}$, it requires superconstant degree $d=\omega(1)$. It is natural to ask whether low error can still be achieved with polynomials of constant degree $d=O(1)$.
As a step in this direction, in this section, we show the existence of a low-error two-source extractor for arbitrary linear min-entropy $k=\delta n$, which is computable by a degree $\leq 4$ polynomial. In contrast, the best known explicit constructions of low-error two-source extractors require min-entropy $k\approx 0.44n$ [Bou05, Lew19]. More precisely, we prove the following.
@@
For every $\delta>0$ there is some $\zeta>0$ such that there exists a $(k=\delta n,\varepsilon=2^{-\zeta n})$-two-source extractor $\mathsf{Ext}:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ that is computable by an $\mathbb{F}_{2}$-polynomial of degree $\leq 4$. Moreover, such a function $\mathsf{Ext}$ can be sampled with probability $0.99$ in time $O(n^{3})$, and $\mathsf{Ext}(x,y)$ can be evaluated in time $O(n^{3})$ for every input $(x,y)$.
We prove Theorem 8 by combining a previously known connection between subspace-evasive sets (over $\mathbb{F}_{2}$) and low-error two-source extractors for linear min-entropy, and a construction of such subspace-evasive sets from degree-$2$ polynomials. This connection was originally conditional on (a weak form of) the approximate duality conjecture [BR15]. However, this conjecture is now a theorem since it is implied by the Polynomial Freiman-Ruzsa theorem, which was recently proved by Gowers, Green, Manners, and Tao [GGMT23].
@@
We say that a function $f:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ has rank at most $r$ if there exist functions $g_{1},g_{2}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{r}$ such that
@@
Equivalently, $f$ has rank at most $r$ if and only if its $2^{n}\times 2^{n}$ truth table has rank at most $r$ over $\mathbb{F}_{2}$.
@@
Ben-Sasson, Lovett, and Ron-Zewi [BLR12] obtained an entropy-error tradeoff for this result, which allows one to drop the min-entropy $k^{\prime}$ to $O(n/\log n)$ while having error $2^{-\Omega(\sqrt{n})}$. For the sake of exposition, we do not use this alternative version, but remark that it can be applied here as well.
@@
A set $S\subseteq\mathbb{F}_{2}^{r}$ is said to be $(\ell,2^{k})$-subspace-evasive if $|S\cap V|<2^{k}$ for any $-\ell$ subspace $V$ of $\mathbb{F}_{2}^{r}$. An injective function $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{r}$ is $(\ell,2^{k})$-subspace-evasive if $h(\mathbb{F}_{2}^{n})$ is $(\ell,2^{k})$-subspace-evasive.
@@
The following lemma relates the subspace-evasiveness of $h$ to the dispersion and rank of $\mathsf{Ext}_{h}$.
@@
Given an injective $(\ell=r/2,2^{k})$-subspace-evasive function $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{r}$, it follows that $\mathsf{Ext}_{h}$ is a $k$-two-source disperser of rank at most $r$.
@@
Fix any two sets $A,B\subseteq\mathbb{F}_{2}^{n}$ of size $2^{k}$ each. We have that $\dim h(A),\dim h(B)\geq r/2+1$, since $h$ is $(r/2,2^{k})$-subspace-evasive. In particular, this means that $\dim h(A)+\dim h(B)>r+1$. By Lemma 17, we conclude that $\mathsf{Ext}_{h}(A,B)=\langle h(A),h(B)\rangle=\{0,1\}$. ∎
@@
For every $C,\gamma,\rho>0$ there exists $\eta>0$ such that the following holds for all $n$: Let $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{r}$ be an injective $(\ell=r/2,2^{k}=2^{\rho n})$-subspace-evasive function with $r\leq Cn$. Then, $\mathsf{Ext}_{h}(x,y)=\langle h(x),h(y)\rangle$ is a $(k^{\prime}=(\rho+\gamma)n,\varepsilon=2^{-\eta n})$-two-source extractor. Moreover, if $h$ is explicit, then so is $\mathsf{Ext}_{h}$.
@@
With Theorem 10 in mind, it remains to construct an $(\ell,2^{k})$-subspace-evasive function $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n+r}$ with $r=O(n)$, $k=o(n)$, and $\ell=\frac{n+r}{2}$. Furthermore, we would like to construct $h$ in a way that $\mathsf{Ext}_{h}$ is computable by a low-degree $\mathbb{F}_{2}$-polynomial. Towards this end, we consider the function
@@
where $f_{1},\dots,f_{r}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ are random degree $d$ polynomials.
@@
Let $k=10dn^{1/d}$ and $r=11n$. Then, with probability at least $1-2^{-3n^{2}}$ over the sampling of $f_{1},\dots,f_{r}$ we have that $h$ is an injective $(\ell=\frac{n+r}{2},2^{k})$-subspace-evasive function.
Sudakov and Tomon [ST23] also consider subspace-evasive sets induced by collections of multivariate polynomials, but they analyze instead the setting where the field is large and both the dimension and intersection size are small. Next, we need the following lemma.
Lemma 19 ([BHL12, Lemma 1.4] and [KS05, Theorem 1.5], adapted).
@@
With this lemma in hand, we are ready to prove Theorem 11.
@@
Set $r=11n$. Sample $r$ random degree $2$ polynomials $f_{1},\dots,f_{r}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ and define $h:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{n+r}$ as $h(x)=(x_{1},\dots,x_{n},f_{1}(x),\dots,f_{r}(x))$. Then, with probability at least $1-2^{-3n^{2}}$ over the sampling of $f_{1},\dots,f_{r}$ the following holds: For any $\delta>0$ there exists $c>0$ such that the function $\mathsf{Ext}_{h}:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ given by
@@
6.2 Improved impossibility results for dispersing from polynomial and variety sources
@@
We show that the same construction we used to obtain low-degree subspace-evasive sets above also works for sumset-evasiveness, if we use our new Lemma 13 instead of Lemma 19. To keep things simple, we focus here on even degree $d$, and note that it is straightforward to extend our results to hold for all degrees $d$, at a slight loss in parameters (as was done for our fully general result on sumset extractors in Section 5.2).
@@
There exist constants $C,c>0$ such that the following holds for all positive integers $k\geq t\geq 200d$ with $d$ even and $\binom{t/100}{\leq d/2}\geq 4d^{2}(2e)^{d}$. Let $f_{1},\dots,f_{r}:\mathbb{F}_{2}^{k}\to\mathbb{F}_{2}$ be independent random degree $d$ polynomials with
@@
which is at most $2^{-r\tau^{\prime}/2}$ provided that $r\geq\frac{4\tau k}{\tau^{\prime}-4\tau}=\frac{4k}{(2e)^{-d}d^{-2}\tau-4}$, the result follows by noting that our hypotheses implies $(2e)^{-d}d^{-2}\tau-4\geq(2e)^{-d}d^{-2}\tau/2$, so it suffices to choose any $r\geq 8d^{2}(2e)^{d}k/\tau$. ∎
@@
for a sufficiently large constant $C>0$. The third equality uses the fact that $k\leq n$ and $\binom{a}{\leq b}\geq\binom{a}{b}\geq\left(\frac{a}{b}\right)^{b}$. Using this and setting $t=\alpha\log n$ for a sufficiently small constant $\alpha>0$ yields the following corollary of Theorem 12, which states that there exist polynomial NOBF sources [^19] Following [CGG24, Definition 4], a degree $d$ polynomial NOBF source $\mathbf{Z}\sim\mathbb{F}_{2}^{n}$ with min-entropy $k$ is a degree $d$ polynomial source with the following structure: There exists a set $G\subseteq[n]$ of $k$ good coordinates which are independent and uniformly distributed, and each coordinate outside $G$ is computed by a degree $\leq d$ polynomial of the coordinates in $G$. Polynomial NOBF sources are a special case of both polynomial sources and variety sources. with extremely high min-entropy that avoid even very tiny sumsets.
@@
For any constant $c>0$ there exists a constant $C>0$ such that the following holds for all positive integers $n\geq C$ and even degrees $d$. There exists a degree $d$ polynomial NOBF source $\mathbf{Z}\sim\mathbb{F}_{2}^{n}$ with min-entropy $H_{\infty}(\mathbf{Z})\geq n-n\left(\frac{Cd}{\log n}\right)^{d/2}$ such that $\mathsf{supp}(\mathbf{Z})$ does not contain any sumset $A+B$ with $|A|,|B|\geq n^{c}$.
Corollary 8 generalizes [CGG24, Theorem 5.11], which worked for $d=2$ only, to all even degrees $d$, and it has a simpler proof. It also yields the following corollary, extending the impossibility result from [CGG24, Section 5.3] on using sumset dispersers to disperse from polynomial NOBF sources in a black-box manner to larger degrees $d>2$.
Corollary 9 (Sumset dispersers cannot disperse from high-entropy polynomial NOBF sources and variety sources, any even degree $d$).
Fix any even degree $d\in\mathbb{N}$. Then, there exists a constant $C>0$ such that for any integer $n\geq C$ the following holds. Sumset dispersers cannot be used to disperse from degree $d$ polynomial NOBF sources over $\mathbb{F}_{2}^{n}$ with min-entropy $k=n-n\left(\frac{Cd}{\log n}\right)^{d/2}$ in a black-box manner. In particular, this also means that sumset dispersers cannot be used to disperse from degree $d$ variety sources with the same min-entropy $k$ in a black-box manner.
@@
In order to be able to conclude that $k^{\prime}$-sumset disperser is also a disperser for another class of sources $\mathcal{C}$ using only its black-box property (i.e., that it is not constant on any sumset $A+B$ with $|A|,|B|\geq 2^{k^{\prime}}$), one needs that for every $\mathbf{X}\in\mathcal{C}$, $\mathsf{supp}(\mathbf{X})$ contains a sumset $A+B$ with $|A|,|B|\geq 2^{k^{\prime}}$. Since we can take $k^{\prime}=\alpha\log n$ for an arbitrarily small constant $\alpha>0$ and there are no sumset dispersers for such $k^{\prime}$ (because every function is constant on an $\Omega(\log n)$-dimensional subspace, and affine sources are sumset sources), the result follows. To see the “In particular” part, it suffices to note that degree $d$ polynomial NOBF sources with min-entropy $k$ are also degree $d$ variety sources with min-entropy $k$ [CGG24, Claim 1]. ∎